OpenStudy (jkbo):

Which equation is in slope-intercept form and represents a line with slope 3 through the point (9, –4)?

OpenStudy (anonymous):

do you have answer choices ?

OpenStudy (anonymous):

there is a formula to use... y - y1 = m(x - x1) slope (m) = 3 (9,-4) -- y1 = -4 and x1 = 9 Now sub all that into the formula Once you get that equation, you can then put it in slope intercept form which is y = mx + b.